OpenStudy (anonymous):

Simplify (3/x + 2/x+2) / (3/x+2 - 2/x)

OpenStudy (anonymous):

\[\frac{ 3 }{ x} + \frac{ 2 }{ x+2 } \div \frac{ 3 }{ x+2 }-\frac{ 2 }{ x }\]

OpenStudy (anonymous):

So I've tried to cancel things out, flipping things around, etc but still can't get the right answer

OpenStudy (anonymous):

A. \[\frac{ -11 }{ 4 }\] B. -1 C. \[\frac{ 5x+6 }{ x-4 }\] D. \[\frac{ 5x + 2 }{ x+2 }\]

OpenStudy (anonymous):

@ganeshie8 you free rn man?

OpenStudy (anonymous):

@dan815 hii are you free rn? Don't wanna bother or anything, just wondering if you could help.

OpenStudy (anonymous):

Cross multiply and then combine like terms.\[\frac{ 3(x+2)+2x }{ 3x-2(x+2) }\]\[\frac{ 3x+6+2x }{ 3x-2x+4 }\]\[\frac{ 5x+6 }{ x-4}\]

OpenStudy (anonymous):

Sorry, +4 in step 2 should be -4.

OpenStudy (anonymous):

whoaaa wait that's all you gotta do? flvs made it seem so much harder

OpenStudy (anonymous):

wait so what would the answer be

OpenStudy (anonymous):

Yep, that's all! It's not as scary as it looks. The answer would be C, then.

OpenStudy (anonymous):

ahh ok thank u very much!! have a nice day xx

OpenStudy (anonymous):

You too. :)